OpenStudy (mejia941):

Seven times a number minus the number is -48. Find the number.

OpenStudy (kkutie7):

you need to start by writing this out in an equation what exactly is this question telling you?

OpenStudy (kkutie7):

lets make x our unknown number: \[7*x-x=-48\] here we have 7 times a number then subtracting that number is -48

jhonyy9 (jhonyy9):

mejia do you can continue it ?

jhonyy9 (jhonyy9):

7x-x = ?

OpenStudy (mejia941):

these are the answers the number is -8 the number is 8 the number is 6

jhonyy9 (jhonyy9):

ok but i like help you solve this exercise step by step - please collaborate

jhonyy9 (jhonyy9):

so there above wrote 7x -x = -48

OpenStudy (mejia941):

thank you

jhonyy9 (jhonyy9):

you have got the right answer ?

jhonyy9 (jhonyy9):

x = ?